Fermi National Laboratory
 

Accelerator Update for July 14, 2006 - July 17, 2006

 

Friday July 14

The day shift began with Operations monitoring store 4828 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, with the NuMI horn cooling work continuing, and with MiniBooNE taking beam.

At 8:46 AM, a Booster RF station (BRF9) tripped off for the first of many times in seven minutes. Operators bypassed the station and called an expert. BRF12 & 15 also started tripping off. Operations reduced the intensity.

A MI expert began a study period at 9 AM.

Operations established beam to Meson MTest at 9:24 AM.

At 9:25 AM, BRF9 tripped off. It reset.

At 9:57 AM, BRF10, 12, & 17 tripped off. The stations all reset.

Operations began preparing the Recycler at 10:34 AM to stash antiprotons.

At 10:43 PM, all the Booster RF stations, except BRF7, tripped off. The West Anode Power Supply also tripped off. A Booster expert reported that the low level RF curves had been corrupted. He fixed the curves and beam resumed at 11:03 AM.

The stash ended and Pbar resumed stacking at 11:29 AM.

MTest resumed taking beam at 11:57 PM.

FESS and LCW technicians reported at 1:52 PM that they had replaced the inlet strainer for the A1 pond. They then headed to CDF to do the same thing.

At 2:31 PM, a Linac RF station (LRF5) tripped off. An expert adjusted the cavity frequency.

FESS and LCW technicians re-cleaned the A1 pond strainer at 3:49 PM.

The Recycler began preparing to stash at 5:28 PM.

At 5:45 PM, Operations reported Recycler correction elements were tripping. Operators contacted an expert who said to put the stash on hold and he would investigate.

The Recycler expert reported at 6:17 PM that he replaced a control card and the bulk power supply, which resolved the problem.

The recycler stash resumed at 7:10 PM.

At 7:25 PM, BRF4 tripped off for the first of four times in five minutes. It reset each time.

At 7:47 PM, Operations reported a problem with 400MeV chopper. Operations called in experts.

The stash ended at Pbar resumed stacking at 7:59 PM.

At 8:35 PM, the H- Source tripped off. It reset okay.

Booster experts reported at 9:18 PM that a multiwire was stuck in the Linac. Operators prepared the area for an access. The experts moved the wire out of the beam path and came out of access at 9:40 PM.

Pbar resumed stacking and MiniBooNE resumed taking beam at 10:12 PM.

Saturday July 15

The midnight shift began with Operations monitoring store 4828 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

The Recycler began preparing to stash at 1:27 AM. The stash ended.

The Recycler began preparing to stash at 4:52 AM. The stash ended and Pbar resumed stacking at 5:18 AM.

At 7:10 AM, an Accumulator RF station (ARF1) tripped off. It reset okay.

The Recycler began preparing to stash at 7:17 AM. The stash ended.

At 7:39 AM, Operations reported that BRF19 had tripped off several times in the last 30 minutes.

Around 8:30 AM, Operations terminated store 4828. Then they put the TeV into shot setup.

At 9:46 AM, Operations reported that Booster had a coalescing problem. They contacted an expert who helped them make adjustments.

At 10:46 AM, Operations reported serious Booster beam position problems. Operators continued to work on the problem.

At 10:49 AM, Operations reported that the MI sequencer would fail every time a shot log command was given and the TeV sequencer would fail on a “COPY_SA” command.

Operations established store 4829 at 11:29 AM with an initial luminosity 135.77E30.

At 11:36 AM, Operations reported that store 4829 aborted. There was no quench. Later, operators discovered that the LCW valve at TeV sector A2 was closed. An expert came in and manually opened it to 50%.

Operations established beam to MTest at 11:52 AM.

Operations began preparing the CDF and D0 collision halls for access around 12:03 PM.

Operations reported at 12:13 PM that the Booster had a kicker timing (MK08) problem.

The Recycler began preparing to stash at 12:54 PM.

At 1:22 PM, Operations reported that the first stash was lost due to a power supply trip.

The stash ended and MTest resumed taking beam.

At 3:32 PM, Operations reported that a MI power supply (MI-60U) tripped off and the VCB opened. Operations called in an expert who had the problem resolved by 4:42 PM. Pbar resumed stacking.

Operations put the TeV through a dry squeeze at 4:55 PM.

At 5:04 PM, BRF8 tripped off for the first of three times in 11 minutes. It reset each time.

At 5:15 PM, Operations reported that a TeV sector A1 separator had sparked twice.

The Recycler began preparing to stash at 6:31 PM. The stash ended and Pbar resumed stacking around 7:05 PM.

Operations reported that TeV experts were conducting studies at 7:56 PM.

The Recycler began preparing to stash at 10:11 PM.

TeV studies ended at 10:24 PM.

At 10:44 PM, Operations reported that the sequencer failed during the stash. The stash ended sometime around 11:24 PM.

Operations put the TeV into shot setup around 11:43 PM.

Sunday July 16

The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

Operations established store 4835 at 1:28 AM with an initial luminosity of 56.65E30.

At 2:02 AM, Operations reported the beam drifting in the AP1 transfer line.

The Recycler began preparing to stash at 3:40 AM. The transfers were very inefficient. Operators investigated and believed that H926 was not making it to its proper current.

At 7:52 AM, Operations reported that Pbar’s PLC02 was down as was LRF5. LRF5 was needed to tune up MI. Operations decided to prepare the Recycler to stash. The stash ended around 8:25 PM. This transfer had a worse efficiency than the last one due to problems with H901 and no MI tune up.

Operations prepared Pbar at 8:34 AM for an access to replace the PLC.

The Pbar access ended and Pbar resumed stacking at 10:45 PM.

The Recycler began preparing to stash at 3:09 PM. At 3:33 PM, Operations reported that D:H926 and D:H916 were tripping off. The stash was called off at 3:46 PM. Operations called in a power supply experts.

Pbar resumed stacking at 3:59 PM.

The Recycler stash resumed at 5:02 PM.

The stash ended and Pbar resumed stacking at 5:42 PM.

At 7:16 PM, Operations reported that BRF17 had tripped off several times within the last two hours.

At 8:09 PM, Operations reported that BRF17 tripped off for the first of four times in twenty minutes. It reset each time.

MTest resumed taking beam at 8:29 PM.

At 9:03 PM, a Debuncher power supply (D:IB) tripped off on a ground fault indication. An operator called in an expert.

At 9:10 PM, the TeV suffered a sector E1 quench. Store 4835 lost. Later, operators discovered that the quench was due to a separator spark. Experts were called in.

The Recycler prepared to stash at 9:17 PM. At 9:29 PM, a Linac quadrupole bulk power supply (Q5BULK) tripped off and then at 9:38 PM, H926 caused a problem during the reverse proton tune up.

At 10:02 PM, the power supply expert reported that D:IB would need to be opened up.

At 10:04 PM, Q5BULK tripped off again and would not reset. Operations called in an expert.

At 10:29 PM, BRF18 tripped off and would not reset. Operators called in experts who had the station fixed shortly before midnight.

Monday July 17

The midnight shift began with Operations monitoring the stash, with the TeV recovering from a quench, with an expert investigating Q5BULK problems, and with MiniBooNE taking beam.

Operations put the TeV through a dry squeeze at 12:49 AM.

The power supply expert working on D:IB reported at 1:20 AM that it was back in service. He replaced two resister banks.

Operations put the TeV into shot setup at 2:48 AM.

Operations established store 4838 at 4:43 AM with an initial luminosity of 79.19E30.

The Recycler prepared to stash at 6:47 PM.

At 7:09 AM, BRF7 tripped off on a group B overload and its modulator would not come on. BRF17 was also off due to no gap envelope. An expert investigated.

The Plans for Monday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/17/2006    email Fermilab